OpenStudy (anonymous):

2y + 8x = 11, what is the value of 3(2y + 8x)?

OpenStudy (anonymous):

33

OpenStudy (anonymous):

how do you work through it though

OpenStudy (anonymous):

multiply 3 and 11. Look at the given relationship

OpenStudy (anonymous):

why would you just multiply them?

OpenStudy (anonymous):

Well since 2y+8x=11 then 3(2y+8x)=3(11)

OpenStudy (anonymous):

I still am confused